OpenStudy (anonymous):

Explain the relationship between -1 power and the original fraction.? Select several fractions between 0 and 1. find the value of each fraction after it is raised to the -1 power.

OpenStudy (ybarrap):

$$ \left (\cfrac{1}{x}\right )^{-1}=\cfrac{1^{-1}}{x^{-1}}=\cfrac{x^1}{1^1}=\cfrac{x}{1}=x $$ So, $$ \left (\cfrac{1}{x}\right )^{-1} $$ Brings the numerator to the denominator and the denominator to the numerator. That is, it "inverts" the faction -- turns it upside down, so-to-speak.

OpenStudy (ybarrap):

For example, $$ \left (\cfrac{1}{2} \right )^{-1}=2 $$